# sources.bib

### v1

@@ 1. Introduction
First introduced in late 2006 by Ginzburg [GinzburgCY] as a way to “transplant conventional Calabi–Yau geometry” into the non-commutative setting, Calabi–Yau algebras are now a mainstay of noncommutative algebra, and occupy a space in many a noncommutative geometer’s consciousness that is difficult to overstate. Often thought of as a good candidate for noncommutative polynomial rings in the $\mathbb{N}$-graded case, the study of $\mathbb{N}$-graded twisted Calabi–Yau algebras encompasses that of Artin–Schelter regular algebras [cytoolkit], and in global dimension $3$ is closely related to path algebras of quivers arising from superpotentials [BOCKLANDT200814]. Work of Reyes and Rogalski in [cytoolkit] has classified locally finite-dimensional $\mathbb{N}$-graded twisted Calabi–Yau algebras in small dimensions, and recent work of Gaddis et al. in [GADDIS202486] classifies certain quivers supporting $\mathbb{N}$-graded twisted Calabi–Yau algebras, extending work done by Rogalski and Rogalski in [GADDIS2021106645]. Inspired by this interest in the $\mathbb{N}$-graded case, our initial motivation is to ask what other types of gradings are supported by these twisted Calabi–Yau algebras.
@@
Let us begin by recalling some of the preliminary definitions. Let $A$ be a $k$-algebra, and define the enveloping algebra of $A$ as $A^{e}\coloneqq A\otimes A^{\operatorname{op}}$, where $A^{\operatorname{op}}$ is the opposite algebra of $A$. As we are assuming that all bimodules are $k$-central, we know that given an $(R,S)$-bimodule $M$, we can view it as a left or right $R\otimes S^{\operatorname{op}}$-module via the action
$(r\otimes s^{\operatorname{op}})\cdot m=r\cdot m\cdot s=m\cdot(s\otimes r^{\operatorname{op}})\text{,}$
where $r\in R,s\in S$, and $m\in M$. In particular, we have a correspondence between left $A^{e}$-modules and $(A,A)$-bimodules. The algebra $A$ is said to be homologically smooth if it has a finite-length resolution of finitely generated projective $A^{e}$-modules; that is, $A$ is perfect as a left $A^{e}$-module. If $\nu$ is an algebra automorphism of $A$, then we let ${}^{1}A^{\nu}$ denote the bimodule $A$ with normal left action, but right action twisted by $\nu$. That is, $a\cdot b\cdot c\coloneqq ab\nu(c)$ for all $a,b,c\in A$. Let us first recall the definition of a twisted Calabi–Yau algebra, before any grading enters the picture.
@@
$\operatorname{Ext}^{i}_{A^{e}}(A,A^{e})\cong\begin{cases}0,\quad&i\neq d\\ U,\quad&i=d\end{cases}$
The use of the term twisted is meant to emphasize that $U$ may not be isomorphic to $A$. If the above isomorphism should hold with $U\cong A$, then we say that $A$ is (untwisted) Calabi–Yau. If the isomorphism were to hold with $U\cong{}^{1}A^{\nu}$ for $\nu$ an algebra automorphism of $A$, then we say that $A$ is twisted Calabi–Yau with Nakayama automorphism $\nu$. We say that the graded $k$-algebra $A=\bigoplus_{n\in\mathbb{N}}A_{\sigma}$ is $\mathbb{N}$-graded twisted Calabi–Yau of dimension $d$ if $A$ is perfect when viewed as an object in the category of $\mathbb{N}$-graded left $A^{e}$-modules, and the above isomorphism holds when viewed as a morphism in the same graded category. The more general $G$-graded situation is defined as follows.
@@
Let $A$ be a $G$-graded $k$-algebra and $U$ a $G$-graded invertible $(A,A)$-bimodule. We say that $A$ is $G$-graded twisted Calabi–Yau of dimension $d$ if $A$ is graded homologically smooth and we have the following graded isomorphism of $(A,A)$-bimodules:
$\operatorname{\underline{EXT}}_{A^{e}}^{i}(A,A^{e})\cong\begin{cases}0,\ i\neq d\\ U,\ i=d\end{cases}$
As we would hope for any good definition, we observe that the familiar $\mathbb{N}$-graded flavor is a special case of this new definition by recalling that any $\mathbb{N}$-grading induces a $\mathbb{Z}$-grading with trivial negative degree terms. Reyes and Rogalski prove in Theorem 4.2 of [cytoolkit] that an $\mathbb{N}$-graded algebra that happens to be twisted Calabi–Yau is the same as an $\mathbb{N}$-graded Calabi–Yau algebra. We find the same holds true in the group-graded setting, as well.
@@
With this theorem in hand, we now have an easy way of locating examples of group-graded twisted Calabi–Yau algebras from known examples: any algebra that is simultaneously $G$-graded and twisted Calabi–Yau is G-graded twisted Calabi–Yau. We initially would like to survey some of the examples that can occur. We have already seen that the Weyl algebras provide examples of $\mathbb{Z}$-graded Calabi–Yau algebras, and will find further examples arising from generalized Weyl algebras, as well as universal enveloping algebras of finite dimensional Lie algebras.
In Section 4 of this paper, we demonstrate another way of generating $G$-graded Calabi–Yau algebras from known $\mathbb{N}$-graded examples, namely, via localization at left and right denominator subsets.
@@
Let $A$ a $k$-algebra and $S$ a left and right denominator subset of $A$. If $A$ is twisted Calabi–Yau of dimension $d$, then $S^{-1}A$ is twisted Calabi–Yau of dimension $d$.
@@ Acknowledgements
This work constitutes a portion of the author’s PhD thesis. She is gratefully advised by Manuel Reyes, and thanks him for his feedback, encouragement, and proofreading of this paper.
@@ 2. Homologically smooth algebras and invertible bimodules: The group-graded flavors
Let us first agree on a few conventions surrounding graded rings and their module categories. Let $G$ denote an abelian group with identity $e$, and whose operation we write multiplicatively. For a ring $R$ to be $G$-graded means that $R$ has a decomposition into additive subgroups $R=\bigoplus_{\sigma\in R}R_{\sigma}$ such that $R_{\tau}R_{\sigma}\subseteq R_{\tau\sigma}$ for all $\tau,\sigma\in G$. Given a $G$-graded ring $R$, by $\mathbf{{}_{R}Mod}$ we mean the usual category of left $R$-modules, and by $\mathbf{G{\text{-}}gr_{R}Mod}$ we mean the category of $G$-graded left $R$-modules. Objects in this graded category are $G$-graded left $R$-modules, and morphisms are the degree-preserving morphisms of $\mathbf{{}_{R}Mod}$. When $G$ is clear from context, we will often write $\mathbf{gr{\text{-}}R}$. Unless otherwise specified, “module” by itself is understood to mean a left module.
As a matter of notation, let $\operatorname{Hom}_{R}(-,-)$ denote homomorphism groups of left modules, and $\operatorname{Hom}_{R^{\operatorname{op}}}(-,-)$ that of right modules. Further, we require that homomorphisms of $R$-modules act opposite that of scalars. Given $M,N$ objects in $\mathbf{gr{\text{-}}R}$ and $f\in\operatorname{Hom}_{R}(M,N)$, we say that $f$ is a graded morphism of degree $\tau\in G$, if $f(M_{\sigma})\subseteq N_{\sigma\tau}$ for all $\sigma\in G$, and we write $f\in\operatorname{Hom}_{R}(M,N)_{\tau}$. Analogously for the right, we say that $f\in\operatorname{Hom}_{R^{\operatorname{op}}}(M,N)$ is a graded morphism of degree $\tau,\tau\in G$, if $f(M_{\sigma})\subseteq\tau\sigma$ for all $\sigma\in G$. We define the “total” Hom-group as follows:
$\operatorname{\underline{HOM}}(M,N)\coloneqq\bigoplus_{\tau\in G}\operatorname{Hom}_{R}(M,N)_{\tau}\subseteq\operatorname{Hom}_{R}(M,N)$
@@ 2.1. Homologically smooth algebras
Define the enveloping algebra of a $k$-algebra $A$ as $A^{e}:=A\otimes_{k}A^{\operatorname{op}}$ and recall that $A$ is said to be homologically smooth if it has a finite length resolution of finitely generated projectives in the category of $A^{e}$-modules; that is, $A$ is perfect as a left $A^{e}$-module. The graded version of this definition is a straightforward extension: if $A$ is $G$-graded, then we consider such a resolution in the $G$-graded $A^{e}$-module category. In developing graded versions of well-loved definitions, a common theme becomes the question of when the graded and ungraded versions are equivalent. As an example of such nuances, consider the fact that a projective object in $\mathbf{gr{\text{-}}R}$ is the same as a graded and projective object in $\mathbf{{}_{R}Mod}$, but the same is not the case if we are to replace “projective” with “injective”, or perhaps even more surprisingly “free” (cf. [sacred2, p.21]). Luckily for us, the graded and ungraded definitions of homologically smooth interact nicely. The following result is an adaptation of [cytoolkit, Lemma 2.4], where it was proven in the $\mathbb{N}$-graded case.
@@
$P_{n}\oplus Q_{n-1}\cdots\cong\ker(\varepsilon_{n-1})\oplus P_{n-1}\oplus\cdots;$
@@ 2.2. Invertible bimodules
As was done for homologically smooth algebras, we now wish to extend results regarding invertible bimodules to their $G$-graded counterparts. Let $A$ a $G$-graded $k$-algebra. In the ungraded world, an $(A,A)$-bimodule $U$ is said to be invertible if there exists an $(A,A)$-bimodule $V$ such that $U\otimes_{A}V\cong V\otimes_{A}U\cong A$ as $(A,A)$-bimodules. To define the $G$-graded version, we need first make sense of the graded tensor product, which we denote by $\underline{\otimes}$. We say that the degree $\gamma$ component of $U\underline{\otimes}_{A}V$ is the vector subspace spanned by all pure tensors of the form $u_{\tau}\otimes v_{\sigma}$ such that $\tau\sigma=\gamma$. That is, $(U\underline{\otimes}_{A}V)_{\gamma}=\bigoplus_{\sigma\tau=\gamma}U_{\sigma}\otimes V_{\tau}$. Now, if $U$ is an $(A,A)$-bimodule that carries a $G$-grading, then we say that $U$ is $G$-graded invertible if there exists $V$ a $G$-graded $(A,A)$-bimodule such that $U\underline{\otimes}_{A}V\cong V\underline{\otimes}_{A}U\cong A$ as graded $(A,A)$-bimodules.
@@
Let $A$ be a $G$-graded $k$-algebra and $U$ a $G$-graded $(A,A)$-bimodule. Then $U$ is invertible if and only if it is $G$-graded invertible.
@@
Let $A$ be a $G$-graded $k$-algebra and $U$ a $G$-graded $(A,A)$-bimodule. Suppose that $U$ is invertible. Then, $U\otimes_{A}-:\mathbf{A{\text{-}}Mod}\to\mathbf{A{\text{-}}Mod}$ provides an autoequivalence of the category of left $A$ modules, and as such $U$ is finitely generated and projective on the left and right, as well as a generator for both $\mathbf{A{\text{-}}Mod}$ and $\mathbf{Mod\text{-}A}$. Finite generation of $U$ as a left module gives equality of the total Hom-group $\operatorname{\underline{HOM}}_{A}(U,A)$ and $\operatorname{Hom}_{A}(U,A)$ on the level of sets, and the map which sends a graded morphism to its ungraded counterpart provides an isomorphism of abelian groups. We similarly have $\operatorname{\underline{HOM}}_{A^{\operatorname{op}}}(U,A)\cong\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)$ as abelian groups.
Now, standard results of Morita theory (c.f. [sacred3, Lemma 18.17]) assert that
@@
Let $\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)\otimes_{A}U=\bigoplus_{\sigma\in G}M_{\sigma}$ where $M_{\sigma}$ is the additive subgroup generated by all elements of the form $f\otimes x$ where $f\in\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)$ has degree $\tau$ and $x\in U$ has degree $\gamma$ such that $\tau\gamma=\sigma$. Consider $f$ and $x$ with these degrees, and that $\Phi(f\otimes x)=f(x)\in A_{\tau\sigma=\gamma}$, so in particular $\Phi(M_{\gamma})\subseteq A_{\gamma}$ for all $\gamma\in G$. Hence, $\Phi$ provides the following graded isomorphism of $A^{e}$-modules:
$\Phi:\operatorname{\underline{HOM}}_{A^{\operatorname{op}}}(U,A)\underline{\otimes}_{A}U\xrightarrow{\sim}A$

### v2

@@ 1. Introduction
First introduced in late 2006 by Ginzburg [GinzburgCY] as a way to “transplant conventional Calabi–Yau geometry” into the non-commutative setting, Calabi–Yau algebras are now a mainstay of noncommutative algebra, and occupy a space in many a noncommutative geometer’s consciousness that is difficult to overstate. Often thought of as a good candidate for noncommutative polynomial rings in the $\mathbb{N}$-graded case, the study of $\mathbb{N}$-graded twisted Calabi–Yau algebras encompasses that of Artin–Schelter regular algebras [cytoolkit], and in global dimension $3$ is closely related to path algebras of quivers arising from superpotentials [BOCKLANDT200814]. Work of Reyes and Rogalski in [cytoolkit] has classified locally finite-dimensional $\mathbb{N}$-graded twisted Calabi–Yau algebras in small dimensions, and recent work of Gaddis et al. in [GADDIS202486] classifies certain quivers supporting $\mathbb{N}$-graded twisted Calabi–Yau algebras, extending work done by Gaddis and Rogalski in [GADDIS2021106645]. Inspired by this interest in the $\mathbb{N}$-graded case, our initial motivation is to ask what other types of gradings are supported by these twisted Calabi–Yau algebras.
@@
Let us begin by recalling some of the preliminary definitions. Let $A$ be a $k$-algebra, and define the enveloping algebra of $A$ as $A^{e}\coloneqq A\otimes A^{\operatorname{op}}$, where $A^{\operatorname{op}}$ is the opposite algebra of $A$. As we are assuming that all bimodules are $k$-central, we know that given an $(A,A)$-bimodule $M$, we can view it as a left or right $A^{e}$-module via the action
$(a\otimes b^{\operatorname{op}})\cdot m=a\cdot m\cdot b=m\cdot(b\otimes a^{\operatorname{op}})\text{,}$
where $a,b\in A$, and $m\in M$. In particular, we have a correspondence between left $A^{e}$-modules and $(A,A)$-bimodules. The algebra $A$ is said to be homologically smooth if it has a finite-length resolution of finitely generated projective $A^{e}$-modules; that is, $A$ is perfect as a left $A^{e}$-module. If $\nu$ is an algebra automorphism of $A$, then we let ${}^{1}A^{\nu}$ denote the bimodule $A$ with normal left action, but right action twisted by $\nu$. That is, $a\cdot b\cdot c\coloneqq ab\nu(c)$ for all $a,b,c\in A$. Let us first recall the definition of a twisted Calabi–Yau algebra, before any grading enters the picture.
@@
$\operatorname{Ext}^{i}_{A^{e}}(A,A^{e})\cong\begin{cases}0,\quad&i\neq d;\\ U,\quad&i=d.\end{cases}$
The use of the term twisted is meant to emphasize that $U$ may not be isomorphic to $A$. If the above isomorphism should hold with $U\cong A$, then we say that $A$ is (untwisted) Calabi–Yau. If the isomorphism were to hold with $U\cong{}^{1}A^{\nu}$ for $\nu$ an algebra automorphism of $A$, then we say that $A$ is twisted Calabi–Yau with Nakayama automorphism $\nu$. We say that the graded $k$-algebra $A=\bigoplus_{\sigma\in\mathbb{N}}A_{\sigma}$ is $\mathbb{N}$-graded twisted Calabi–Yau of dimension $d$ if $A$ is perfect when viewed as an object in the category of $\mathbb{N}$-graded left $A^{e}$-modules, and the above isomorphism holds when viewed as a morphism in the same graded category. The more general $G$-graded situation is defined as follows.
@@
Let $A$ be a $G$-graded $k$-algebra, and let $U$ be a $G$-graded invertible $(A,A)$-bimodule. We say that $A$ is $G$-graded twisted Calabi–Yau of dimension $d$ if $A$ is graded homologically smooth and we have the following graded isomorphism of $(A,A)$-bimodules:
$\operatorname{\underline{EXT}}_{A^{e}}^{i}(A,A^{e})\cong\begin{cases}0,\ i\neq d;\\ U,\ i=d.\end{cases}$
As we would hope for any good definition, we observe that the familiar $\mathbb{N}$-graded flavor is a special case of this new definition. Indeed, any $\mathbb{N}$-graded algebra $A=\bigoplus_{n\in\mathbb{N}}A_{n}$ can be realized as a $\mathbb{Z}$-graded algebra by setting $A_{n}=0$ for $n<0$. Reyes and Rogalski prove in Theorem 4.2 of [cytoolkit] that an $\mathbb{N}$-graded algebra that happens to be twisted Calabi–Yau is the same as an $\mathbb{N}$-graded twisted Calabi–Yau algebra. We find the same holds true in the group-graded setting, as well.
@@
With this theorem in hand, we now have an easy way of locating examples of group-graded twisted Calabi–Yau algebras from known examples: any algebra that is simultaneously $G$-graded and twisted Calabi–Yau is G-graded twisted Calabi–Yau. We initially would like to survey some of the examples that can occur. In addition to the Weyl algebras, we present further examples of $\mathbb{Z}$-graded Calabi–Yau algebras including examples arising from generalized Weyl algebras, as well as universal enveloping algebras of finite dimensional Lie algebras.
In Section 4 of this paper, we use localization at left and right denominator subsets to generate $G$-graded Calabi–Yau algebras from known $\mathbb{N}$-graded examples.
@@
Let $A$ be a $k$-algebra, and let $S$ be a left and right denominator subset of $A$. If $A$ is twisted Calabi–Yau of dimension $d$, then $S^{-1}A$ is twisted Calabi–Yau of dimension $d$.
@@ Acknowledgements
This work constitutes a portion of the author’s PhD thesis. She is advised by Manuel Reyes, and thanks him for his feedback, encouragement, and proofreading of this paper. She also thanks the referee for their careful reading of the paper and providing suggestions which have improved its quality and clarity.
@@ 2. Homologically smooth algebras and invertible bimodules: The group-graded flavors
Let us first agree on a few conventions surrounding graded rings and their module categories. Let $G$ denote an abelian group with identity $e$, and whose operation we write multiplicatively. For a ring $R$ to be $G$-graded means that $R$ has a decomposition into additive subgroups $R=\bigoplus_{\sigma\in G}R_{\sigma}$ such that $R_{\tau}R_{\sigma}\subseteq R_{\tau\sigma}$ for all $\tau,\sigma\in G$. Unless otherwise specified, “module” by itself is understood to mean a left module. Let $R$ be a $G$-graded ring. For an $R$-module $M$ to be $G$-graded means that $M$ has a decomposition into additive subgroups $M=\bigoplus_{\sigma\in G}M_{\sigma}$ such that $R_{\tau}M_{\sigma}\subseteq M_{\tau\sigma}$ for all $\tau,\sigma\in G$. Given a $G$-graded ring $R$, by $\mathbf{{}_{R}Mod}$ we mean the usual category of left $R$-modules, and by $\mathbf{G{\text{-}}gr_{R}Mod}$ we mean the category of $G$-graded left $R$-modules. Objects in this graded category are $G$-graded left $R$-modules, and morphisms are the degree-preserving morphisms of $\mathbf{{}_{R}Mod}$. When $G$ is clear from context, we will often write $\mathbf{gr{\text{-}}R}$.
Let $\operatorname{Hom}_{R}(-,-)$ denote homomorphism groups of left modules, and let $\operatorname{Hom}_{R^{\operatorname{op}}}(-,-)$ denote those of right modules. Further, we require that homomorphisms of $R$-modules act opposite that of scalars. Given $M,N$ objects in $\mathbf{gr{\text{-}}R}$ and $f\in\operatorname{Hom}_{R}(M,N)$, we say that $f$ is a graded morphism of degree $\tau\in G$, if $f(M_{\sigma})\subseteq N_{\sigma\tau}$ for all $\sigma\in G$, and we write $f\in\operatorname{Hom}_{R}(M,N)_{\tau}$. Analogously for the right, we say that $f\in\operatorname{Hom}_{R^{\operatorname{op}}}(M,N)$ is a graded morphism of degree $\tau\in G$, if $f(M_{\sigma})\subseteq\tau\sigma$ for all $\sigma\in G$. We define the “total” Hom-group as follows:
$\operatorname{\underline{HOM}}(M,N)\coloneqq\bigoplus_{\tau\in G}\operatorname{Hom}_{R}(M,N)_{\tau}\subseteq\operatorname{Hom}_{R}(M,N).$
@@ 2.1. Homologically smooth algebras
Define the enveloping algebra of a $k$-algebra $A$ as $A^{e}:=A\otimes_{k}A^{\operatorname{op}}$ and recall that $A$ is said to be homologically smooth if it has a finite length resolution of finitely generated projectives in the category of $A^{e}$-modules; that is, $A$ is perfect as a left $A^{e}$-module. The graded version of this definition is a straightforward extension: if $A$ is $G$-graded, then we consider such a resolution in the $G$-graded $A^{e}$-module category. In developing graded versions of definitions, a common theme becomes the question of when the graded and ungraded versions are equivalent. As an example of such nuances, consider the fact that a projective object in $\mathbf{gr{\text{-}}R}$ is the same as a graded and projective object in $\mathbf{{}_{R}Mod}$, but the same is not the case if we are to replace “projective” with “injective”, or perhaps even more surprisingly “free” (cf. [sacred2, p.21]). However, the graded and ungraded definitions of homologically smooth interact nicely. The following result is an adaptation of [cytoolkit, Lemma 2.4], where it was proven in the $\mathbb{N}$-graded case.
@@
$P_{n}\oplus Q_{n-1}\cdots\cong\ker(\varepsilon_{n-1})\oplus P_{n-1}\oplus\cdots.$
@@ 2.2. Invertible bimodules
As was done for homologically smooth algebras, we now wish to extend results regarding invertible bimodules to their $G$-graded counterparts. Let $A$ a $G$-graded $k$-algebra. In the ungraded world, an $(A,A)$-bimodule $U$ is said to be invertible if there exists an $(A,A)$-bimodule $V$ such that $U\otimes_{A}V\cong V\otimes_{A}U\cong A$ as $(A,A)$-bimodules. To define the $G$-graded version, we need first make sense of the graded tensor product, which we denote by $\underline{\otimes}$. We say that the degree $\gamma$ component of $U\underline{\otimes}_{A}V$ is the vector subspace spanned by all pure tensors of the form $u_{\tau}\otimes v_{\sigma}$ such that $\tau\sigma=\gamma$. That is, $(U\underline{\otimes}_{A}V)_{\gamma}=\bigoplus_{\sigma\in G}U_{\gamma\sigma^{-1}}\otimes_{A}V_{\sigma}$. Now, if $U$ is an $(A,A)$-bimodule that carries a $G$-grading, then we say that $U$ is $G$-graded invertible if there exists $V$ a $G$-graded $(A,A)$-bimodule such that $U\underline{\otimes}_{A}V\cong V\underline{\otimes}_{A}U\cong A$ as graded $(A,A)$-bimodules.
@@
Let $A$ be a $G$-graded $k$-algebra, and let $U$ be a $G$-graded $(A,A)$-bimodule. Then $U$ is invertible if and only if it is $G$-graded invertible.
@@
Let $A$ be a $G$-graded $k$-algebra, and let $U$ be a $G$-graded $(A,A)$-bimodule. Suppose that $U$ is invertible. Then, $U\otimes_{A}-:\mathbf{A{\text{-}}Mod}\to\mathbf{A{\text{-}}Mod}$ provides an autoequivalence of the category of left $A$ modules, and as such $U$ is finitely generated and projective on the left and right, as well as a generator for both $\mathbf{A{\text{-}}Mod}$ and $\mathbf{Mod\text{-}A}$. Finite generation of $U$ as a left module gives equality of the total Hom-group $\operatorname{\underline{HOM}}_{A}(U,A)$ and $\operatorname{Hom}_{A}(U,A)$ on the level of sets, and the map which sends a graded morphism to its ungraded counterpart provides an isomorphism of abelian groups. We similarly have $\operatorname{\underline{HOM}}_{A^{\operatorname{op}}}(U,A)\cong\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)$ as abelian groups.
Now, standard results of Morita theory (cf. [sacred3, Lemma 18.17]) assert that
@@
Let $\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)\otimes_{A}U=\bigoplus_{\sigma\in G}M_{\sigma}$ where $M_{\sigma}$ is the additive subgroup generated by all elements of the form $f\otimes x$ where $f\in\operatorname{Hom}_{A^{\operatorname{op}}}(U,A)$ has degree $\tau$ and $x\in U$ has degree $\gamma$ such that $\tau\gamma=\sigma$. Consider $f$ and $x$ with these degrees, and that $\Phi(f\otimes x)=f(x)\in A_{\tau\gamma=\sigma}$, so in particular $\Phi(M_{\sigma})\subseteq A_{\sigma}$ for all $\sigma\in G$. Hence, $\Phi$ provides the following graded isomorphism of $A^{e}$-modules:
$\Phi:\operatorname{\underline{HOM}}_{A^{\operatorname{op}}}(U,A)\underline{\otimes}_{A}U\xrightarrow{\sim}A.$